Suppose an economy can be represented by the following table, in which employment is in millions of workers and GDP and AE are expressed in billions of dollars: Employment Real GDP 100 105 110 115 120 125 1200 1300 1400 1500 1600 1700 Aggregate Expenditures 1275 1350 1425 1500 1575 1650 Use the table to answer the following: What is the equilibrium level of GDP? size? GDP. What is the multiplier in this economy? below the economys potential, what is the size of the recessionary expenditure gap? a. b. What kind of expenditure gap exists if full employment is 120 million workers? What is its c. Suppose government spending, taxes, and net exports are all independent of the level of real d. Suppose instead that the economy is producing at equilibrium GDP. If this GDP is $200 billion
Answer #1

a) Equilibrium level of GDP is at the point where - Actual GDP (real GDP) equals Potential GDP (aggregate expenditure)

The level of employment where this happens is at 115, where Real GDP = Aggregate expenditure = $1500 billion

b) At 120 employment, Real GDP (1600) > Potential GDP (1575). Such expenditure gap is called inflationary gap where real GDP > potential GDP. This shows that economy is operating at the GDP more than its full potential GDP.

The size of inflationary gap is - (1600 - 1575) = $25 billion.

c) If the government spending , taxes and net exports are all independent of the level of real GDP, that will mean that level of real GDP will be just dependent on the consumption spending.

The consumption multiplier = 1 / (1-MPC) ,where MPC = Marginal propensity to consume.

Thus multiplier in this economy is - 1/(1-MPC)

d) At equilibrium GDP, if the GDP is $200 billion below the economy's potential , that means that real GDP is (1500 - 200 = 1300) and potential GDP is the same as 1500. This means now,

Real GDP < Potential GDP , (this shows recessionary gap in the economy)

Recessionary gap = Real GDP - Potential GDP

= 1300 - 1500

= - $200 billion